OpenStudy (anonymous):

6y^2-24y=0 help solve with a little explanation plzz :0

OpenStudy (anonymous):

elekt u posted but it disappeared -_-

OpenStudy (anonymous):

Factor out the y \[y^2-24y=\] \[y(6y-24)=0\]

OpenStudy (anonymous):

okay i understand so far

OpenStudy (anonymous):

Ah, I always mess up with writing equations in here. Anyway, so y = 0 or y = 4

OpenStudy (anonymous):

okay im not following how you got the solution from y(6y-24)=0

OpenStudy (anonymous):

You just factor out the y, because it is a common factor in the expression.

OpenStudy (anonymous):

Try mulitplying the expression y(6y-24) to get back 6y^2-24y again

OpenStudy (anonymous):

yes i get that but how do you get to the sol. set (0,4)

OpenStudy (anonymous):

to get the solutions it to equal 0 either y=0 or 6y-24=0 y=4 in that case. Following on from Elekt original message

OpenStudy (anonymous):

Thanks Maclntosh :)

OpenStudy (anonymous):

how do you get y=4?

OpenStudy (anonymous):

You see that, in order for the expression to equal zero, once factored, either y must equal zero or the parantheses must equal zero. Multiplying by zero always gives zero, remember :)

OpenStudy (anonymous):

y=4 you get from solving 6y-24 = 0 like this 6y = 24 y = 24/6 = 4

OpenStudy (anonymous):

Complete solution: \[6y^2-24y= 0 \] \[y(6y-24)=0\] In order for this to happen \[y=0\] or \[6y-24 = 0\] \[6y =24\] \[y = \frac{24}{6} = 4\]

OpenStudy (anonymous):

6y-24=0 <-- thats after we factored out a y, right? so where did the y go?

OpenStudy (anonymous):

i get now, how to get y=4 but theres a lost y :(

OpenStudy (anonymous):

No, the factored y is already accounted for because it must equal 0 in order for the equation to be satisfied. If it is not zero, well then y = 4 because of the above.

OpenStudy (anonymous):

so once its factored out its gone for good? ;p i kinda get it i guess

OpenStudy (anonymous):

No, not at all. We just see that the equation y(6y-24=0 must have either what is outside the parantheses, that is y, equal to zero, or the inside of the parantheses, that is 6y-24, equal to zero for the equation to be satisfied.

OpenStudy (anonymous):

okay now that the y=4 is understood can you help with how you got the y=0?

OpenStudy (anonymous):

ohhh wow.. gotcha

OpenStudy (anonymous):

Thus y must equal either 0 or 4 for this to hold :)

OpenStudy (anonymous):

okayy. this is so confusing and i have a test on this thur! what should i do to practice and learn this betterr?

OpenStudy (anonymous):

Understand it.

OpenStudy (anonymous):

okay! so lets to this prob.. ill tell u what i think i should do..

OpenStudy (anonymous):

x^2=-49

OpenStudy (anonymous):



OpenStudy (anonymous):

This won't work, because something squared is never negative (when you work with real numbers anyway. Square roots can be defined for negative numbers in something called the complex plane of numbers, but that is far above your head at the moment).

OpenStudy (anonymous):

x=-7i or 7i

OpenStudy (anonymous):

We're working on basic algebra here, introducing complex numbers will just confuse things.

OpenStudy (anonymous):

no im suppose to use imaginary numbers!

OpenStudy (anonymous):

Oh, u are?

OpenStudy (anonymous):

no this is from complex ch

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, than I am sorry to assume you were not working with that *flush*

OpenStudy (anonymous):

its okay! so help explain how kavu got his answer \

OpenStudy (anonymous):

x^2=-49 x^2=√(7)^2√(-1) or x^2=√(-7)^2√(-1) x=7i or x=-7i